# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v2

@@ Abstract
This paper presents a minimal operator-only term rewriting system with seven constructors and eight reduction rules. The main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. From strong normalization, a certified normalizer with proven totality and soundness is derived. Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment. Impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication, are established. The work demonstrates fundamental limitations in termination proving for self-referential systems. A conjecture is stated: no relational operator-only TRS can have its full-system termination proved by internally definable methods. Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting. Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures. This structural limitation applies to any TRS expressive enough to encode ordered data. All theorems have been formally verified in a proof assistant. The Lean formalization is available at https://github.com/MosesRahnama/OperatorKO7.
@@ 1 Introduction
Highlights (formalization summary).
@@ 3 The KO7 calculus
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing. R_rec_succ redistributes $s$ to two RHS positions.
Kernel signature (7 constructors).
@@
Tiny example (trace consequences).
@@ 4 Strong normalization
Step vs SafeStep (measure unification).
@@
No-Go for constant bumps on $\kappa$ (generic duplicator).
@@
Duplication stress identity.
@@ 6 Local confluence and Newman (guarded safe relation)
Scope and Guarantees (formalization-accurate).
@@
  • Full Step per-rule decreases (Hybrid). For each kernel rule, there is a per-step decrease witnessed either by KO7’s $\delta/\kappa^{M}/\mu$ lex or by an MPO-leaning $\mu$-first triple. A uniform global aggregator for all of Step is left as future work.
Critical-pair coverage.
@@
$\delta$-guard: definition and decidability.
@@
7 Impossibility results
The formal development establishes the failure of several simpler measure strategies, necessitating the use of DM multiset orders or MPO:
  • Additive bumps fail: No measure of the form $\kappa(t)+k$ strictly decreases on generic duplicating rules.
  • Bare flags fail: A single boolean flag is insufficient to orient rules that lift depth.
  • Polynomial interpretations fail: Any polynomial interpretation involving fixed constants requires external arithmetic axioms to “orient” the rule, violating the operator-only constraint.
Polynomial Impossibility.
@@ 7 Impossibility results
Internally definable measures.
@@ 9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means.
A relational operator-only TRS is, by definition, one capable of ordered computation. The terms are equivalent: “relational” = “capable of ordered computation.” Such systems can express iteration over successors, comparison of structure depths, or any form of sequential counting. The minimal signature for such capability includes a recursor, an operator that applies a step function iteratively across a successor-structured argument. Examples include primitive recursion over natural numbers, fold over lists, and any construct that steps through an ordered index. A TRS that cannot express such iteration (e.g., a flat pattern-matching system with no recursive structure) is not relational in this sense.
The critical property of relational systems is that their recursor redistributes its step argument across recursive calls. This redistribution is not optional: it is what makes ordered computation possible. But it also defeats all additive termination measures.
The structural barrier.
@@
Scope of the conjecture.
The conjecture applies to any TRS expressive enough to encode ordered data (natural numbers, lists, trees with successor structure, or any type supporting iteration). Systems without such structure (purely ground, non-recursive, or lacking a step-iteration pattern) fall outside the conjecture’s scope. The claim is that once a system has enough structure for ordered computation, its full termination escapes internal proof methods.
Evidence.
@@ 10 Formalization structure
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available at https://github.com/MosesRahnama/OperatorKO7.
Acknowledgments
AI has been a great help with this project.
Appendix: Addendum
Lean kernel location
@@ Acknowledgments
Triple-lexicographic measure and duplication handling (DM/MPO)
@@
Aggregation for $\kappa^{M}$ (union, not sum).
@@
Short witness snippets (toy duplication).
@@
Newman scope (guarded safe relation)
@@
rec succ $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
@@
Table 2: Termination Proof Strategy: Per-rule Orientation Summary.
@@
Local-join report (Confluence)
@@
Appendix: Orientation Strategy
@@
Appendix A Related work
This work follows standard expositions in Baader-Nipkow and Terese [2, 11]. Termination under duplication relies on DM multiset orders [5] and (recursive) path orders [6]. Newman’s Lemma is used in its classic ARS form [10]. For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [7, 8, 3]. For fixed-target reachability complexity and modularity results, see [1, 4]; for undecidability at shallow depth in non-linear systems, see [9].
Appendix B Conclusion

### v3

@@ Abstract
This paper presents a minimal operator-only term rewriting system (KO7) designed to investigate the foundations of self-verifying computation. The system consists of seven constructors and eight reduction rules, with no external axioms, no imported logic, and no pre-existing arithmetic.
Standard termination methods fail on this system. We document ten distinct proof strategies that succumb to the system’s structural complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, and unchecked recursion. These methods collapse at the same barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats every additive measure ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). This failure is not a defect of the proof search but a necessary feature of any system capable of sequential computation.
To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit. For this fragment, a novel triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and ordinal ranking) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
We state a fundamental conjecture: No relational operator-only TRS can have its full-system termination proved by internally definable methods. We define a “relational” TRS as one capable of internalized sequential computation. This property is shared by Turing machines (via transition functions), lambda calculus (via fixed-point combinators), and recursive functions. All such systems require a recursor to define the “next” step; this recursor inherently creates the duplication barrier.
This result reveals a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), this barrier appears at a more primitive layer: in operator systems possessing only order and recursion, before arithmetic emerges. The system’s inability to prove its own termination from within is a structural limitation of ordered computation itself, paralleling the halting problem but manifesting at the substrate of operator definitions.
The complete Lean 4 formalization ( 7,000 lines) is available at https://github.com/MosesRahnama/OperatorKO7.
@@ 1 Introduction
Highlights (formalization summary)
@@ 3 The KO7 calculus
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing.
@@
Kernel signature (7 constructors)
@@
Tiny example (trace consequences)
@@ 4 Strong normalization
Step vs SafeStep (measure unification)
@@
No-Go for constant bumps on $\kappa$ (generic duplicator)
@@
Duplication stress identity
@@ 6 Local confluence and Newman (guarded safe relation)
Scope and Guarantees (formalization-accurate)
@@
Critical-pair coverage
@@
Table 2: Critical-pair coverage at the root for SafeStep.
@@
$\delta$-guard: definition and decidability
@@
7 Impossibility results: The Ten Failed Methods
The formal development establishes the failure of ten distinct termination strategies, necessitating the use of the triple-lexicographic measure with Dershowitz-Manna orderings. These failures are not merely heuristic; they are formally witnessed by counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Operational_Incompleteness.lean).
  1. 1.
    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
  2. 2.
    Simple Lexicographic Measures: Pairs like $(\kappa,\mu)$ fail because the primary $\kappa$ component does not strictly decrease on duplicating steps.
  3. 3.
    Polynomial Interpretations: Without external arithmetic coefficients (banned by the operator-only constraint), polynomial measures degrade to additive ones and fail on duplication.
  4. 4.
    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void‘) can toggle the flag from 0 to 1.
  5. 5.
    Pure Ordinal Measures: Shape-blind ordinals fail to distinguish nested $\delta$ structures from their contexts.
  6. 6.
    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like merge_cancel where the result may have equal depth to the input.
  7. 7.
    Naive Multiset Orderings: Standard multiset measures fail without the specific properties of the Dershowitz-Manna order (which allows a single element to be replaced by any number of smaller elements).
  8. 8.
    Local Helper Lemmas: Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
  9. 9.
    Complex Hybrid/Constellation Measures: Attempts to combine measures in ad-hoc ways failed to provide a uniform decrease across all 8 rules.
  10. 10.
    Unchecked Recursion: The raw rec_succ rule itself serves as the ultimate counterexample; without the SafeStep guards, it permits unbounded recursion that no internal measure can bound.
These ten failures collectively demonstrate that the termination barrier is structural: the system’s internal logic cannot capture the necessary invariants without the sophisticated machinery of the triple-lexicographic measure.
Polynomial Impossibility
@@ 7 Impossibility results: The Ten Failed Methods
Internally definable measures
@@ 9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means
A relational operator-only TRS is one capable of sequential computation: any system that can express “do $X$, then do YY” without external conditions. The terms are equivalent: “relational” = “capable of sequential computation.”
9.1 The Relational Nature of KO7 and Recursors
The classification of KO7 as “relational” is not arbitrary; it describes the system’s capability to internalize sequential control flow. In flat rewrite systems (e.g., group axioms), rules rewrite terms to equivalent terms. In a relational TRS like KO7, the rewrite rules define a computational process.
[…]
$\mathrm{rec}\Delta(b,s,\delta(n))\to\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the transition function $\delta$ of a Turing machine or the fixed-point combinator $Y$ in lambda calculus. Both mechanisms allow a system to execute a sequence of operations of unbounded length.
The Recursor Trap
Any system capable of this sequential computation must have a mechanism to carry the “step” logic ($s$) forward into the next iteration. In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7), this redistribution of $s$ defeats all additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A system that internalizes its own control flow in this way cannot prove its own termination using simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
Escaping the Trap: Meta-Layers vs. Internal Proofs
Standard systems escape this trap by stepping outside the system.
  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally. The system doesn’t prove itself; the compiler proves the system.
  • Turing Machines: We do not ask the tape to prove it halts; we analyze the machine from the outside (and finding that we cannot, we discover the Halting Problem).
KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag). The Conjecture posits that for the full, unguarded system, no such internal proof exists.
The structural barrier
@@ 9.1 The Relational Nature of KO7 and Recursors
Scope of the conjecture
The conjecture applies to any operator-only TRS capable of sequential computation: any system that can express “do $X$, then do YY” using only rewrite rules, without external axioms, imported arithmetic, or borrowed logic. Systems without such capability (purely ground, non-recursive, or lacking step-iteration) fall outside the conjecture’s scope. The claim is that once a system can compute sequentially without external conditions, its full termination escapes internal proof methods.
Evidence
@@ 10 Formalization structure
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development (Lean 4,  7,000 LOC; Apache-2.0 licensed) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
11 Conclusion
Appendix A Addendum
A.1 Lean kernel location
@@
A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
@@
Aggregation for $\kappa^{M}$ (union, not sum)
@@
Short witness snippets (toy duplication)
@@
A.3 Newman scope (guarded safe relation)
@@ A.4 Module map
Table 3: Per-rule decrease witnesses used in the SafeStep termination proof.
@@
rec succ [^2] This row records the per-step decrease used in the SafeStep proof. It does not assert termination of the unrestricted full relation: rec_succ duplicates the step argument $s$ and is the structural source of the full-system termination barrier (cf. the conjecture/evidence discussion). $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
@@
A.6 Local-join report (Confluence)
@@
Appendix B Orientation Strategy